# Spectral Structure of the Mixed Hessian of the Dispersionless Toda $\tau$-Function

### 24 Mar 2026 v1

@@ Abstract.
We analyze the mixed Hessian of the dispersionless Toda $\tau$-function for the $s$-fold symmetric one-harmonic polynomial conformal map. The inverse branch exhibits two distinct thresholds: an analytic threshold $\zeta_{c}$, where the dominant square-root singularity reaches the circle of convergence, and a later geometric threshold $\zeta_{\mathrm{univ}}>\zeta_{c}$, where the map ceases to be univalent. We prove that the first spectral instability occurs already at $\zeta_{c}$. In each symmetry sector, the weighted subcritical realization has exactly one logarithmically diverging eigenvalue, whereas the remaining spectrum stays bounded and, after removal of the singular direction, converges to that of a compact limiting remainder. We further continue the corresponding scalar Gram functions beyond $\zeta_{c}$, showing that they admit a generalized hypergeometric description, a Cauchy–Stieltjes representation, and, for $1\leq p\leq s$, a realization as Weyl functions of bounded Jacobi operators. In particular, these scalar quantities remain finite at $\zeta_{\mathrm{univ}}$. This identifies analytic criticality, rather than loss of univalence, as the first spectral threshold of the Toda Hessian.
@@ 1. Introduction
Planar domains, conformal maps, and harmonic moments play a central role in several closely related subjects, including Laplacian growth, dispersionless integrable hierarchies, Green-function and Bergman-kernel theory, and planar matrix models [17, 24, 16, 13, 9, 25, 23]. A common feature of these theories is that the geometry of a simply connected domain may be encoded by its conformal map and by the harmonic moments of its complement. This makes it natural to ask not only how the domain itself degenerates near criticality, but also how the associated second-order response objects behave. The present paper studies one such object: the mixed Hessian of the dispersionless Toda free energy, equivalently the coefficient matrix of a mixed logarithmic kernel attached to the inverse conformal map.
@@
be the unique exterior conformal map normalized by $f(w)/w\to r$ as $w\to\infty$. The parameter $r>0$ is the conformal radius of $D$, while the remaining coefficients encode the shape of the boundary. Let $w=w(z)$ denote the inverse branch near infinity, normalized by $w(z)=z/r+O(z^{-1})$.
Two bilinear kernels are naturally attached to $f$. The first is the holomorphic Faber–Grunsky kernel $\log\frac{f(w)-f(w^{\prime})}{r(w-w^{\prime})}$ whose coefficients control univalence through the classical Grunsky inequalities [5, 19]. By contrast, the present paper is concerned with the mixed holomorphic–antiholomorphic logarithmic kernel
@@
valid for $|z|$ and $|z^{\prime}|$ sufficiently large. It is obtained by pulling back the canonical mixed logarithmic kernel of the exterior disk under the inverse conformal map. The coefficients $H_{mn}$ are the main object of the present paper.
@@
$H_{mn}=r^{m+n}\frac{\partial^{2}\mathcal{F}}{\partial t_{m}\,\partial\bar{t}_{n}}.$
Thus $(H_{mn})$ is the scale-invariant mixed Hessian of the dispersionless free energy.
The same coefficients may be interpreted in several equivalent languages. In Laplacian growth and Hele–Shaw dynamics, $(H_{mn})$ records the second-order coupling between moment deformations. In the dispersionless Toda hierarchy, it is the matrix of mixed second derivatives of $\log\tau$. In Bergman-kernel language, one has, with our normalization,
@@
so the same data may be recovered from the exterior Bergman kernel in conformal coordinates [3, 11]. In the normal matrix model, $\mathcal{F}$ is the planar free energy, or equivalently the logarithmic energy of the equilibrium measure on $D$ [25, 23], and $(H_{mn})$ may be viewed as a mixed susceptibility matrix with respect to moment deformations.
The main contribution of the paper is to determine the spectral structure of $(H_{mn})$ near criticality in a family for which the inverse map can be controlled explicitly. The question is which part of the spectrum becomes singular near criticality, and whether that transition is governed by analytic criticality of the inverse branch or by the later geometric loss of univalence. This question is genuinely different from the classical holomorphic Grunsky theory, whose operator bounds are tied directly to univalence and quasiconformal extension [19, 5]. In the mixed sector studied here, the first instability is triggered already at the analytic threshold and concentrates into one logarithmically diverging stiff mode per symmetry sector, that is, an eigen-direction whose eigenvalue diverges logarithmically, while the remaining sectorial spectrum stays bounded. The resulting separation between analytic and geometric criticality is invisible in a purely holomorphic framework.
Relation to prior work. The present problem should also be distinguished from spectral and determinantal questions for the classical Grunsky operator and matrix, including Fredholm-eigenvalue problems for quasicircles and recent asymptotics for truncated Grunsky operators [21, 22, 12]. Those works concern the holomorphic Grunsky data, or Fredholm determinants built from them. Here the basic object is instead the mixed Hessian $(H_{mn})$, equivalently the mixed logarithmic kernel (1.1); the relevant transition is therefore not the saturation of a bounded holomorphic operator norm or a truncated determinant asymptotic, but a logarithmic rank-one spike in the mixed sector at the earlier analytic threshold $\zeta_{c}$.
With this motivation in hand, we now restrict to a family in which the analysis can be carried out explicitly. We study the $s$-fold symmetric one-harmonic family
@@
which is the simplest family in which the analytic singularity of the inverse map and the geometric loss of univalence separate. Maps of the form (1.3) are standard in Laplacian growth and related free-boundary problems [9]; they already exhibit nontrivial geometric transitions, including loss of univalence and, for $s\geq 3$, boundary cusps at the geometric threshold. Because they depend on a single dimensionless shape parameter $\zeta:=a/r>0$, they provide a minimal and fully explicit model in which the spectral behavior of the mixed Hessian can be analyzed in detail. All rigorous results proved below concern the family (1.3); the corresponding extension to general polynomial conformal maps, under an isolated dominant square-root orbit hypothesis, is developed in the companion paper [2].
Writing $x:=r/z$ and $w(z)^{-1}=x\,U(x;\zeta)$, one finds that the inverse branch is governed by the algebraic equation
@@
at which the dominant square-root singularity of the inverse branch reaches the unit circle in the $x$-plane. The second is the geometric threshold
@@
at which the conformal map ceases to be univalent and the boundary develops its first geometric singularity: for $s\geq 3$, semicubical cusps, and for $s=2$, the degenerate Joukowski segment. The problem is to determine which of these thresholds governs the first spectral instability of the mixed Hessian.
Our answer is that the relevant threshold is $\zeta_{c}$. The $s$-fold symmetry of (1.3) implies $H_{mn}=0$ unless $m\equiv n\pmod{s}$, so the Hessian decomposes into $s$ independent blocks. These blocks admit an exact positive Gram factorization in terms of Raney coefficients generated by (1.4). After passing to a suitable weighted realization, each symmetry block develops one logarithmically diverging eigenvalue, while the remaining sectorial spectrum stays bounded. Equivalently, there is one stiff mode per symmetry sector, so at most $s$ logarithmically diverging eigenvalues globally. The first spectral transition is therefore an analytic phenomenon of the inverse branch rather than a manifestation of geometric non-univalence.
A second theme of the paper concerns the intermediate regime
$\zeta_{c}<\zeta<\zeta_{\mathrm{univ}},$
where the conformal map remains univalent although the weighted compact operator realization used in the subcritical regime is no longer available. In this region the block-operator picture breaks down, but its scalar Gram building blocks remain well defined and continue analytically. More precisely, the block Gram weights in the subcritical operator theory are obtained from generating functions built from the squared Raney coefficients, and these scalar functions still admit analytic continuation beyond $\zeta_{c}$. We identify these continued functions as generalized hypergeometric objects, derive their resonant branch-point expansion, prove a Cauchy–Stieltjes representation, and, in the positive range $1\leq p\leq s$, obtain a Jacobi-matrix realization. Thus the continuation of these scalar Gram data captures the same analytic branch-point mechanism that produces the subcritical logarithmic instability, even after the weighted compact realization has ceased to exist. In particular, the singularity at $\zeta_{c}$ is not the onset of geometric breakdown: the two lateral boundary values of the continued scalar data remain finite at the univalence threshold.
@@ 1.1. Main results
The results of the paper fall into three groups. Theorem A gives the operator-theoretic description of the analytic threshold $\zeta_{c}$: for a fixed weight parameter $\beta>0$, each symmetry block splits into one logarithmically stiff direction and a bounded soft remainder. Theorem B identifies the scalar Gram data that survive beyond the weighted compact regime and continues them across the cut as generalized hypergeometric/Stieltjes objects. Finally, Proposition C shows that this analytic transition occurs strictly before the geometric loss of univalence at $\zeta_{\mathrm{univ}}$.
Fix a symmetry sector $q\in\{1,\dots,s\}$ and a weight parameter $\beta>0$. Let $H^{(q)}$ denote the intrinsic coefficient matrix of the $q$-th Hessian block, and let $\widetilde{G}^{(q)}(\zeta)$ be the corresponding weighted block Gram operator. The spectral analysis in this paper is carried out for these weighted realizations on the fixed Hilbert space $\ell^{2}(\mathbb{N}_{0})$, where $\mathbb{N}_{0}:=\{0,1,2,\dots\}$. The weighting is only an auxiliary Hilbert-space device: it does not alter the coefficient data and does not remove the critical instability. The weighted Gram and weighted Hessian blocks have the same nonzero spectrum, so the spectral conclusions may be stated equivalently for $\widetilde{G}^{(q)}(\zeta)$ or $\widetilde{H}^{(q)}(\zeta)$.
Theorem (A: Rank-one logarithmic instability).
@@
All remaining sectorial eigenvalues stay bounded; more precisely, after removing the spike direction, each fixed soft eigenvalue converges to the corresponding eigenvalue of the compressed compact limit described in Proposition 4.8.
Theorem A is proved by combining the square-root singularity analysis of the inverse branch with the blockwise Gram representation of the Hessian. The logarithmic term is first isolated at the level of matrix entries, and the weighted realization for fixed $\beta>0$ then yields a compact operator framework in which the rank-one singular part separates from the bounded soft sector.
Remark 1.1 (Contrast with the Grunsky operator).
The classical Grunsky operator belongs to the holomorphic sector and is governed by the sharp bound $\|B\|\leq 1$, which is equivalent to univalence of the underlying map [19]. Accordingly, in the present one-parameter family, the holomorphic-sector transition is governed by the geometric threshold $\zeta_{\mathrm{univ}}$. By contrast, the mixed Hessian studied here transitions earlier, at the analytic threshold $\zeta_{c}$, and the nature of the transition is different: instead of the saturation of a bounded operator norm, one eigenvalue diverges logarithmically while the remaining spectrum stays bounded. This distinction between the holomorphic and mixed sectors is genuinely operator-theoretic and is not visible at the level of individual matrix entries.
@@
(1.5) $\mathcal{G}_{p}(u):=\sum_{m\geq 0}R_{s,p}(m)^{2}\,u^{m},\qquad u=\zeta^{2}.$
Below $\zeta_{c}$, the scalar Gram weights entering the block operators are obtained from these functions by explicit Euler-type differential operators. The importance of $\mathcal{G}_{p}$ is therefore that they encode, in scalar form, the same branch-point mechanism that drives the logarithmic spectral instability. These functions continue analytically beyond the radius of convergence $u=\zeta_{c}^{2}$, even though the weighted compact operator picture no longer persists there.
Theorem (B: Analytic continuation of the scalar Gram data).
@@
with $A$ and $B$ analytic near $\zeta_{c}^{2}$. The scalar Gram weights are recovered from $\mathcal{G}_{p}$ by explicit Euler-type differential operators; these remove the quadratic prefactor and produce the logarithmic divergence at $u=\zeta_{c}^{2}$.
Theorem B is obtained from the explicit coefficient ratio of $R_{s,p}(m)^{2}$, which identifies $\mathcal{G}_{p}$ with a generalized hypergeometric function. The local expansion (1.6) is then derived by Frobenius analysis at the regular singular point $u=\zeta_{c}^{2}$. Moreover, the boundary values across the cut define a discontinuity density, and $\mathcal{G}_{p}$ admits a Cauchy–Stieltjes representation. In the positive range $1\leq p\leq s$, the corresponding Stieltjes measure is positive, which yields a Jacobi-matrix realization of $\mathcal{G}_{p}$ as a Weyl $m$-function.
@@
Proposition (C: Separation of thresholds).
@@
Hence the logarithmic spectral instability of Theorem Theorem occurs while the conformal map is still univalent and the boundary is still smooth. Moreover, the analytically continued scalar data remain finite at $\zeta=\zeta_{\mathrm{univ}}$ (Propositions 6.3 and 6.5).
Thus the singularity of the Hessian at $\zeta_{c}$ is not a consequence of cusp formation. It is already encoded in the branch-point structure of the inverse map and therefore precedes geometric breakdown.

### 16 Apr 2026 v2

@@ Abstract.
We study the mixed Hessian of the dispersionless Toda $\tau$-function for the one-harmonic $s$-fold symmetric conformal map $f(w)=rw+aw^{1-s}$. This Hessian is the susceptibility matrix generated by the inverse conformal map. Our spectral statements are formulated for its weighted symmetry-block realizations on a fixed Hilbert space. In that realization, the first spectral transition occurs at the analytic threshold $\zeta_{c}=(s-1)^{s-1}/s^{s}$, where the dominant square-root singularity of the inverse map reaches the normalization circle, rather than at the geometric threshold $\zeta_{\mathrm{univ}}=1/(s-1)$, where univalence fails. After symmetry decomposition and weighted realization, each block develops exactly one logarithmically diverging eigenvalue as $\zeta\uparrow\zeta_{c}$, while the remaining spectrum stays bounded and converges to a compact limit. The instability is therefore rank one in every symmetry sector of the weighted block theory.
We then continue the scalar Gram generating functions beyond $\zeta_{c}$. They are generalized hypergeometric functions on the slit plane $\mathbb{C}\setminus[\zeta_{c}^{2},\infty)$, their branch-point expansion contains the logarithmic term responsible for the divergence, and in the range $1\leq p\leq s$ they admit Cauchy–Stieltjes and Jacobi-matrix realizations. In particular, the continued scalar quantities remain regular at $\zeta_{\mathrm{univ}}$, so the analytic spectral transition strictly precedes the geometric breakdown of univalence.
@@ 1. Introduction
The dispersionless Toda hierarchy associates to a planar domain a $\tau$-function whose mixed second derivatives with respect to harmonic moments define a natural susceptibility matrix. For simply connected domains, this mixed Hessian is generated by the inverse conformal map and therefore encodes analytic information about the domain in a form adapted to integrable-hierarchy methods [16, 25, 15, 12, 23, 8, 26, 24]. In the present paper we study this Hessian for the explicit one-harmonic $s$-fold symmetric family $f(w)=rw+aw^{1-s}$, which provides a nontrivial solvable case at the interface of conformal mapping, dispersionless integrability, and spectral analysis.
Our main observation is that the first spectral instability of the weighted symmetry-block realization occurs strictly before the conformal map ceases to be univalent. More precisely, we identify an analytic threshold $\zeta_{c}=(s-1)^{s-1}/s^{s}$, at which the dominant square-root singularity of the inverse map reaches the normalization circle, and we prove that this threshold is strictly smaller than the geometric threshold $\zeta_{\mathrm{univ}}=1/(s-1)$ for loss of univalence. The spectral statement is formulated for the weighted block realizations on a fixed Hilbert space introduced below: in that realization, each symmetry block develops exactly one logarithmically diverging eigenvalue as $\zeta\uparrow\zeta_{c}$, while all remaining eigenvalues stay bounded. Thus the onset of instability is rank one in every symmetry sector.
@@
be the unique exterior conformal map normalized by $f(w)/w\to r$ as $w\to\infty$. Writing $w=w(z)$ for the inverse map near infinity, one obtains the mixed logarithmic kernel
@@
valid for $|z|$ and $|z^{\prime}|$ sufficiently large. It is obtained by pulling back the canonical mixed logarithmic kernel of the exterior disk under the inverse conformal map. In the dispersionless Toda framework this is precisely the Wiegmann–Zabrodin mixed kernel identity [25, 12]. The coefficients $H_{mn}$ are the main object of the present paper.
@@
$H_{mn}=r^{-(m+n)}\frac{\partial^{2}\mathcal{F}}{\partial t_{m}\,\partial\bar{t}_{n}}.$
Thus $(H_{mn})$ is the scale-invariant mixed Hessian of the dispersionless free energy. This interpretation is closely connected with Takhtajan’s free-boson approach to the $\tau$-function for smooth Jordan contours [23]. In the genus-zero case, $\log\tau$ is a Kähler potential for a natural Hermitian metric on the space of contours of fixed area, so the mixed derivatives of $\log\tau$ represent the corresponding metric coefficients in harmonic-moment coordinates.
The same coefficients therefore appear naturally in several standard frameworks. In Laplacian growth and Hele–Shaw dynamics, $(H_{mn})$ describes the second-order coupling between moment deformations. In the dispersionless Toda hierarchy, it is the matrix of mixed second derivatives of $\log\tau$. In Bergman-kernel language, one has, with our normalization,
@@
so the same coefficients may be recovered from the exterior Bergman kernel in conformal coordinates [2, 10]. In the normal matrix model, $\mathcal{F}$ is the planar free energy, or equivalently the logarithmic energy of the equilibrium measure on $D$ [26, 24], and $(H_{mn})$ may be viewed as a mixed susceptibility matrix with respect to moment deformations.
The main contribution of the paper is to identify the first singularity seen by the block spectrum of the mixed Hessian and to show that this singularity is analytic rather than geometric. This question is genuinely different from the classical holomorphic Grunsky theory, whose operator bounds are tied directly to univalence and quasiconformal extension [18, 4]. In the mixed sector studied here, the first instability already appears at the analytic threshold and concentrates into one logarithmically diverging stiff mode per symmetry sector, while the remaining sectorial spectrum stays bounded. The resulting separation between analytic and geometric criticality is invisible in a purely holomorphic framework. It should also be distinguished from spectral and determinantal questions for the classical Grunsky operator, including recent asymptotics for truncated holomorphic Grunsky operators [20, 21, 11]: those works concern the holomorphic Grunsky coefficients or Fredholm determinants built from them, whereas here the basic object is the mixed Hessian $(H_{mn})$, and the relevant transition is a logarithmic rank-one spike at the earlier analytic threshold $\zeta_{c}$.
We now restrict to a family in which the analysis is completely explicit. We study the $s$-fold symmetric one-harmonic family
@@
which is the simplest family in which the analytic singularity of the inverse map and the geometric loss of univalence separate. Maps of the form (1.3) are standard in Laplacian growth and related free-boundary problems [8]. They already exhibit nontrivial geometric transitions, including loss of univalence and, for $s\geq 3$, boundary cusps at the geometric threshold. Because they depend on a single dimensionless shape parameter $\zeta:=a/r>0$, they provide a minimal and fully explicit example in which the spectral behavior of the mixed Hessian can be analyzed in detail. The present paper is deliberately restricted to this one-harmonic family. Its purpose is to isolate the threshold mechanism and the resulting rank-one spectral instability in the simplest case where both can be made completely explicit.
Writing $x:=r/z$ and $w(z)^{-1}=x\,U(x;\zeta)$, one finds that the inverse map is governed by the algebraic equation
@@
at which the dominant square-root singularity of the inverse map reaches the unit circle in the $x$-plane. The second is the geometric threshold
@@
at which the conformal map ceases to be univalent and the boundary develops its first geometric singularity: for $s\geq 3$, semicubical cusps, and for $s=2$, the degenerate Joukowski segment. The problem is to determine which of these thresholds governs the first spectral instability seen by the weighted block realizations of the mixed Hessian.
For the weighted block Hessian, the relevant threshold is $\zeta_{c}$. The $s$-fold symmetry of (1.3) implies $H_{mn}=0$ unless $m\equiv n\pmod{s}$, so the Hessian decomposes into $s$ independent blocks. These blocks admit an exact positive Gram factorization in terms of Raney coefficients generated by (1.4). The spectral theorem is proved for the associated weighted block realizations on the fixed Hilbert space $\ell^{2}(\mathbb{N}_{0})$: the coefficient matrix is intrinsic, while the weighting is the method that makes the logarithmic instability visible in a fixed operator framework. In this realization, each symmetry block develops exactly one logarithmically diverging eigenvalue, while the remaining sectorial spectrum stays bounded. Equivalently, there is one stiff mode per symmetry sector, so at most $s$ logarithmically diverging eigenvalues globally. The first spectral transition is therefore an analytic phenomenon of the inverse map rather than a manifestation of geometric non-univalence.
A second part of the paper concerns the intermediate regime, $\zeta_{c}<\zeta<\zeta_{\mathrm{univ}}$, where the conformal map remains univalent although the weighted compact operator realization from the subcritical regime is no longer available. The block-operator framework therefore disappears, but the scalar objects underlying the Gram construction remain meaningful and continue analytically. More precisely, the Gram weights of the subcritical theory are generated by series built from the squared Raney coefficients, and these scalar functions admit analytic continuation beyond $\zeta_{c}$. We identify the continued functions as generalized hypergeometric objects, derive their resonant branch-point expansion, establish a Cauchy–Stieltjes representation, and, for $1\leq p\leq s$, obtain a Jacobi-matrix realization. Thus the continued scalar Gram quantities retain the same analytic branch-point mechanism that underlies the subcritical logarithmic instability, even though the weighted compact realization is no longer available. In particular, the singularity at $\zeta_{c}$ is not a geometric one: the two lateral boundary values of the continued scalar quantities remain finite at the univalence threshold.
@@ 1.1. Main results
The main results address three points. Theorem A gives the operator-theoretic form of the analytic threshold $\zeta_{c}$: for a fixed weight parameter $\beta>0$, each symmetry block splits into one logarithmically stiff direction and a bounded soft remainder. Theorem B identifies the scalar Gram quantities that survive beyond the weighted compact regime and continues them across the cut as generalized hypergeometric and Stieltjes objects. Proposition C then shows that this analytic transition occurs strictly before the geometric loss of univalence at $\zeta_{\mathrm{univ}}$.
Fix a symmetry sector $q\in\{1,\dots,s\}$ and a weight parameter $\beta>0$. Let $H^{(q)}$ denote the intrinsic coefficient matrix of the $q$-th Hessian block, and let $\widetilde{G}^{(q)}(\zeta)$ be the corresponding weighted block Gram operator. The spectral analysis in this paper is carried out for these weighted realizations on the fixed Hilbert space $\ell^{2}(\mathbb{N}_{0})$, where $\mathbb{N}_{0}:=\{0,1,2,\dots\}$. This point is important: the coefficient matrix $H^{(q)}$ is intrinsic, but the spectral theorem concerns the weighted operators $\widetilde{G}^{(q)}(\zeta)$ and equivalently $\widetilde{H}^{(q)}(\zeta)$. The weighting therefore does not modify the coefficients. Rather, it places the problem in a fixed Hilbert-space framework in which the critical rank-one spike can be isolated. Since the weighted Gram and weighted Hessian blocks have the same nonzero spectrum, the spectral conclusions may be stated equivalently for $\widetilde{G}^{(q)}(\zeta)$ or $\widetilde{H}^{(q)}(\zeta)$.
Theorem A (Rank-one logarithmic instability).
@@
All remaining sectorial eigenvalues stay bounded. More precisely, after removing the spike direction, each fixed soft eigenvalue converges to the corresponding eigenvalue of the compressed compact limit described in Proposition 4.8.
Theorem A is proved by combining the square-root singularity analysis of the inverse map with the blockwise Gram representation of the Hessian. The logarithmic term is first isolated at the level of matrix entries, and the weighted realization for fixed $\beta>0$ then yields a compact operator framework in which the rank-one singular part separates cleanly from the bounded soft sector.
@@
(1.5) $\mathcal{G}_{p}(u):=\sum_{m\geq 0}R_{s,p}(m)^{2}\,u^{m}.$
For the subcritical conformal family one has the specialization $u=\zeta^{2}$. Below $\zeta_{c}$, the Gram weights entering the block operators are obtained from these functions by explicit Euler-type differential operators. The importance of $\mathcal{G}_{p}$ is therefore that, in scalar form, they capture the same branch-point mechanism that drives the logarithmic spectral instability. These functions continue analytically beyond the radius of convergence $u=\zeta_{c}^{2}$, even though the weighted compact operator picture no longer persists there.
Theorem B (Analytic continuation of the generating function).
@@
with $A$ and $B$ analytic near $\zeta_{c}^{2}$. The scalar Gram weights are recovered from $\mathcal{G}_{p}$ by explicit Euler-type differential operators. These remove the quadratic prefactor and produce the logarithmic divergence at $u=\zeta_{c}^{2}$.
Theorem B is obtained from the explicit coefficient ratio of $R_{s,p}(m)^{2}$, which identifies $\mathcal{G}_{p}$ with a generalized hypergeometric function. The local expansion (1.6) is then derived by Frobenius analysis at the regular singular point $u=\zeta_{c}^{2}$. Moreover, the boundary values across the cut define a discontinuity density, and $\mathcal{G}_{p}$ admits a Cauchy–Stieltjes representation. In the positive range $1\leq p\leq s$, the corresponding Stieltjes measure is positive by the Hausdorff-moment theorem of Liu–Pego [13, 14]. As a consequence, $\mathcal{G}_{p}$ admits a Jacobi-matrix realization as a Weyl $m$-function.
@@
Proposition C (Separation of thresholds).
@@
Hence the logarithmic spectral instability of Theorem A in the weighted block realization occurs while the conformal map is still univalent and the boundary is still smooth. Moreover, the analytically continued scalar quantities remain finite at $\zeta=\zeta_{\mathrm{univ}}$ (Propositions 6.3 and 6.5).
Thus the singularity seen in the weighted mixed-Hessian block theory at $\zeta_{c}$ is not a consequence of cusp formation. It is already encoded in the branch-point structure of the inverse map and therefore precedes geometric breakdown.